# Understanding the Potential of FPGA-Based Spatial Acceleration for Large Language Model Inference

### 23 Dec 2023 v1

@@ 7. Discussion
In this section, we delve into several unanswered questions and open challenges.
AI-Optimized FPGAs. In §4, we demonstrate the potential of leveraging FPGAs with specialized compute engines to accelerate LLMs. Although AIEs and tensor blocks provide massive compute power (Xilinx, 2022a; Langhammer et al., 2021), the memory layout and bandwidth requirements remain undiscovered. Future FPGAs for AI workloads should provide enough memory bandwidth and efficient on-chip interconnect to facilitate the local data movements in a spatial architecture. Moreover, these specialized hardware blocks usually adopt a unique programming model with custom compilation flows. It is still an open question whether existing practices for programming those hardware blocks enable efficient execution of Transformer models.
Timing Closure on Multi-Die FPGAs. We encounter timing problems in partitioning and scaling our design in §5.2. In general, it is hard to adequately explore the design space of multi-die partitioning and scaling. There are automated frameworks (Guo et al., 2021; Khatti et al., 2023) to generate floorplanning constraints, but they are currently not expressive enough to capture the various data movement schemes (e.g., residual connection, multi-head splitting) within Transformer models. We hope similar tools for Transformers could be derived from our analytical framework to speed up the design closure.
Heterogeneous Deployment. Nowadays, data centers are increasingly heterogeneous, with CPUs, GPUs, and FPGAs available at scale (Caulfield et al., 2016; Chung et al., 2018; Narayanan et al., 2021b). Therefore, it is possible to leverage the advantages of different hardware to accelerate Transformer models. For example, GPUs are good for the GPT prefill stage due to their high compute power; FPGAs can achieve low-latency decode stage with customized spatial architecture. The key challenge is to build a distributed system that efficiently manages hundreds of heterogeneous devices. We hope our analysis on resource constraints, latency, and scaling could assist future deployment and evaluation of LLMs in a heterogeneous and distributed environment.
@@ 8. Related Work
FPGA-Based Transformer Accelerators. Most of the prior works on hardware accelerators leverage temporal or overlay architecture with one FPGA (Liu et al., 2021; Li et al., 2020a; Qi et al., 2021a; Hur et al., 2023; Khan et al., 2021; Peng et al., 2021; Li et al., 2021; Qi et al., 2021b). Their performance usually suffers from frequent data movements of intermediate results. DFX (Hong et al., 2022) explores using multiple FPGAs to accelerate GPT2 inference, but it is still an overlay design. Some research has delved into software-hardware co-design to optimize the attention kernel (Zhang et al., 2021b). These endeavors often lack in-depth analysis on resource utilization and cannot be easily generalized to other kernels.
Quantization on LLMs. Initial investigations (Xiao et al., 2023; Dettmers et al., 2022; Yao et al., 2022) demonstrate lossless 8-bit quantization for LLMs. Subsequent studies (Frantar et al., 2022; Lin et al., 2023; Yao et al., 2022; Yuan et al., 2023; Kim et al., 2023a; Zhao et al., 2023) keep lowering the bit width; the latest advancements reveal that 2-bit (Chee et al., 2023) and even 1-bit (binary) quantization (Zhang et al., 2023) are adequate for an accurate LLM. While these approaches offer valuable insights, our focus remains orthogonal to quantization, as we illustrate optimization techniques and provide high-performance building blocks for deploying quantized LLMs on FPGAs.
HLS Kernel Libraries. Despite the existence of kernel libraries for accelerating Transformer models on GPUs (Dao et al., 2022; Wolf et al., 2019; Lefaudeux et al., 2022), the hardware domain has seen only a handful of initiatives in this regard. AMD provides Vitis HLS library (Xilinx, 2022d, e) that only has basic kernel-level examples without comprehensive designs tailored for Transformer models. Some frameworks map deep learning frameworks to FPGAs (Zhang et al., 2018; Fahim et al., 2021; Umuroglu et al., 2017; Blott et al., 2018; Zhang et al., 2020), but can only handle small CNN designs and do not cater to LLMs. More recent tools allow hardware design using Python (Lai et al., 2019; Xiang et al., 2022; Huang et al., 2021; Ye et al., 2022), but are still general-purpose and require hardware engineers to construct and optimize kernels from scratch. Our work provides a Transformer kernel library designed for dataflow implementations and demonstrates their composability in constructing high-performance hardware accelerators.

### 07 Apr 2024 v2

@@ 7. Discussion
In the previous sections, we provide details of the analytical framework and prove that it can achieve high accuracy compared to the latency of actual implementation. However, our framework may have limitations when analyzing the overlay designs or compressed models with sparsity, which requires changes in the resource and latency estimation. In this section, we will delve into several unanswered questions and open challenges.
AI-Optimized FPGAs. In §4, we demonstrate the potential of leveraging FPGAs with specialized compute engines to accelerate LLMs. Although AIEs and tensor blocks provide massive compute power (xilinxAIE, ; stratix10, ), the memory layout and bandwidth requirements remain undiscovered. Future FPGAs for AI workloads should provide enough memory bandwidth and efficient on-chip interconnect to facilitate the local data movements in a spatial architecture. Moreover, these specialized hardware blocks usually adopt a unique programming model with custom compilation flows. It is still an open question whether existing practices for programming those hardware blocks enable efficient execution of Transformer models.
Timing Closure on Multi-Die FPGAs. We encounter timing problems in partitioning and scaling our design in §5.2. In general, it is hard to adequately explore the design space of multi-die partitioning and scaling. There are automated frameworks (guo2021autobridge, ; moazin2023pasta, ) to generate floorplanning constraints, but they are currently not expressive enough to capture the various data movement schemes (e.g., residual connection, multi-head splitting) within Transformer models. We hope similar tools for Transformers could be derived from our analytical framework to speed up the design closure.
Heterogeneous Deployment. Nowadays, data centers are increasingly heterogeneous, with CPUs, GPUs, and FPGAs available at scale (caulfield2016microsoft, ; chung2018brainwave, ; narayanan2021megatronv2, ). Therefore, it is possible to leverage the advantages of different hardware to accelerate Transformer models. For example, GPUs are good for the GPT prefill stage due to their high compute power; FPGAs can achieve low-latency decode stage with customized spatial architecture. The key challenge is to build a distributed system that efficiently manages hundreds of heterogeneous devices. We hope our analysis on resource constraints, latency, and scaling could assist future deployment and evaluation of LLMs in a heterogeneous and distributed environment.
@@ 8. Related Work
FPGA-Based Transformer Accelerators. Most of the prior works on hardware accelerators leverage temporal or overlay architecture with one FPGA (liu2021fqbert, ; li2020ftrans, ; qi2021iccad, ; hur2023flexrun, ; khan2021npe, ; peng2021cbbp, ; li2021nmta, ; qi2021bmc, ). Their performance usually suffers from frequent data movements of intermediate results. DFX (hong2022dfx, ) explores using multiple FPGAs to accelerate GPT2 inference, but it is still an overlay design. Some research has delved into software-hardware co-design to optimize the attention kernel (zhang2021attn_codesign, ). These endeavors often lack in-depth analysis on resource utilization and cannot be easily generalized to other kernels.
Quantization on LLMs. Initial investigations (xiao2023smoothquant, ; dettmers2022llmint8, ; yao2022zeroquant, ) demonstrate lossless 8-bit quantization for LLMs. Subsequent studies (frantar2022gptq, ; lin2023awq, ; yao2022zeroquant, ; yuan2023rptq, ; kim2023squeezellm, ; zhao2023atom, ) keep lowering the bit width; the latest advancements reveal that 2-bit (chee2023quip, ) and even 1-bit (binary) quantization (zhang2023binarized, ) are adequate for an accurate LLM. While these approaches offer valuable insights, our focus remains orthogonal to quantization, as we illustrate optimization techniques and provide high-performance building blocks for deploying quantized LLMs on FPGAs.
HLS Kernel Libraries. Despite the existence of kernel libraries for accelerating Transformer models on GPUs (dao2022flashattention, ; wolf2019huggingface, ; xFormers2022, ), the hardware domain has seen only a handful of initiatives in this regard. AMD provides Vitis HLS library (vitis_hls_library, ; vitis_ai, ) that only has basic kernel-level examples without comprehensive designs tailored for Transformer models. TRAC (patrick2022trac, ) attempts to provide an HLS-based Transformer library, but its kernel performance is unpredictable, and it exclusively focuses on the BERT model using a temporal architecture. Some frameworks map deep learning frameworks to FPGAs (zhang2018dnnbuilder, ; fahim2021hls4ml, ; yaman2017finn, ; blott2018finnr, ; zhang2020dnnexplorer, ), but can only handle small CNN designs and do not cater to LLMs. More recent tools allow hardware design using Python (lai2019heterocl, ; xiang2022heteroflow, ; huang2021pylog, ; yehpca2022scalehls, ; debjit2022dac, ), but are still general-purpose and require hardware engineers to construct and optimize kernels from scratch. Our work provides a Transformer kernel library designed for dataflow implementations and demonstrates their composability in constructing high-performance hardware accelerators.
@@ 9. Conclusion
Acknowledgements.
This work was supported in part by ACE, one of the seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA and NSF Awards #2007832, #2019306, and #2118709. We would like to thank anonymous reviewers, Keisuke Kamahori, and Zihao Ye for providing insightful feedback. We also thank Jiajie Li, Jie Liu, and Zhanqiu Hu for their contributions to the initial LLM modeling and benchmarking.